OpenStudy (anonymous):

The point (2, 3) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Make sure to show all work.

OpenStudy (michele_laino):



OpenStudy (anonymous):

So can you explain that? lol

OpenStudy (michele_laino):

it is the representation of your problem

OpenStudy (michele_laino):

now, we can write this: \[\Large OP = \sqrt {{2^2} + {3^2}} = ...\]

OpenStudy (michele_laino):

what is OP?

OpenStudy (michele_laino):

I have applied the Theorem of Pitagora to the triangle OP1P

OpenStudy (anonymous):

the sin tan and cos?

OpenStudy (michele_laino):

yes, we have to compute OP first

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

hint: \[\Large OP = \sqrt {{2^2} + {3^2}} = \sqrt {4 + 9} = ...?\]

OpenStudy (anonymous):

is it just square root of 13

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

for all three

OpenStudy (michele_laino):

no, since we have to apply these formulas: \[\Large \begin{gathered} \cos \theta = \frac{{O{P_1}}}{{OP}} = \frac{2}{{\sqrt {13} }} = ... \hfill \\ \hfill \\ \sin \theta = \frac{{P{P_1}}}{{OP}} = \frac{3}{{\sqrt {13} }} = ... \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

finally, by definition, we can write: \[\Large \tan \theta = \frac{{P{P_1}}}{{O{P_1}}} = ...\]

OpenStudy (anonymous):

3/2?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

sweet thanks for the help!

OpenStudy (michele_laino):

:)